OpenStudy (anonymous):

what is the measures of a 20-gon

OpenStudy (aum):

The sum of all external angles of a n-gon adds to 360 degrees. A 20-gon has 20 external angles and their sum equals 360 degrees. Therefore, each external angle = 360 / 20 = 18 degrees. Each internal angle = 180 - external angle = 180 - 18 = 162 degrees.